OpenStudy (anonymous):

What is the value of the expression? A. 15.5 B. 3.25 C. 3.875 D. 7.75

OpenStudy (anonymous):

im confusing myself with all the numbers

OpenStudy (imstuck):

Start with the 8+2. Then divide 30 by that sum. Then cube that quotient. Then add 4 to that. You will get a certain number. For the next part, take 2 to the 4th, then divide that product by 4. You will get another number. Divide the first "certain number" by that "another number" and you will get your answer! Let me know if you still get caught up in all the numbers. Ok?

OpenStudy (imstuck):

TY for the medals, guys!

OpenStudy (anonymous):

Always start from the inner parentheses. Then you work your way out.

OpenStudy (anonymous):

Ok - 8+2 is 10, 10 divided by 30 is 3. How do you cube that? sorry, I can never understand these

OpenStudy (anonymous):

To cube, you multiply 3*3*3

OpenStudy (anonymous):

3x3x3=27

OpenStudy (imstuck):

Three to the third means 3*3*3. That gives you...........now add 4 to that 27.

OpenStudy (anonymous):

Then add 4.

OpenStudy (anonymous):

31

OpenStudy (anonymous):

Now you can do the other side.

OpenStudy (anonymous):

good

OpenStudy (anonymous):

2x2x2x2=16

OpenStudy (anonymous):

16 divided by 4 is4

OpenStudy (anonymous):

so 31 divided by 4 is...?

OpenStudy (anonymous):

7.75 I believe

OpenStudy (anonymous):

7.75 !

OpenStudy (anonymous):

yay! we did it!

OpenStudy (anonymous):

The exponent on a number says how many times the number multiplies by itself. 3^3=3*3*3 4^5=4*4*4*4*4

OpenStudy (anonymous):

Thank you all. @Dscdago @lacrosseplayer22 @IMStuck

OpenStudy (anonymous):

:) ur welcome

OpenStudy (anonymous):

I'm glad to help :)

OpenStudy (anonymous):

same